OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = 4, an = 2 • an-1

OpenStudy (anonymous):

If you read the formula, it says: if you want to know the term an, you have to double the previous one. So every new term is 2 times the previous term. Sounds easy enough, but... where to start? You have to know at least one term to be able to get the next. No problem! You have a1, which is 4. So now you can easily double that, to find a2, double that to find a3, doubl...............you get it I think!

OpenStudy (anonymous):

its 2*a n-1 though so it wouldnt just be doubling youd have to subtract something im just not sure if its the term they give you or by the number of the term your finding (ie a1,a2,a3 etc.

OpenStudy (blurbendy):

you can find a(2) a(2) = 2 * 1 we substitute 1, because a(2 -1) = a(1) = 1 so a(2) = 2 a(3) = 2 * 2 a(3) =4 a(4) = 2 * 4 a(4) = 8 a(5) = 2* 8 a(5) = 16

OpenStudy (blurbendy):

ziko1995 was right